OpenStudy (anonymous):

the derivitve of ln(1-3x)

OpenStudy (anonymous):

d/dx (log(1-3x)) = 3/3x-1

OpenStudy (anonymous):

ln(1-3x) its derivative 1/(1-3x) * -3=-3/(-3x+1)

OpenStudy (anonymous):

Because ln(1-3x) follows the form ln u, you'll need to use the chain rule. Like Nick said, you multiply d/dx (ln u) by u'. If you take the derivative of u = 1-3x, you get -3. So -3*(1/1-3x) = d/dx[ln (1-3x)]

OpenStudy (anonymous):

...Did you post that in the right place, Kyleigh?